OpenStudy (anonymous):

Fan and Medal The square of a number decreased by 25 is -9. A. Find the x-intercepts. B. Find the vertex of the graph. C. What shape is the graph? Describe what you see.

OpenStudy (dan815):

x^2-25=-9

OpenStudy (dan815):

x^2-25+9=0 solve for x they are the x intercepts

OpenStudy (dan815):

graph the functions y=x^2-25+9

OpenStudy (anonymous):

Im confused with the graphing part.

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@ajanijones